OpenStudy (anonymous):

Permutation or Combination? I'm stuck on word problems .. such as this one: A local vendor is giving away treats. Two baskets contain 12 bags of chips and 10 candy bars. If you randomly select one treat from each basket, what is the probability that both of the treats are candy bars?

OpenStudy (amistre64):

niether P or C?

OpenStudy (amistre64):

does each basket contian 12 and 10 equally?

OpenStudy (anonymous):

it doesn't really say, that's all the word problem says

OpenStudy (amistre64):

22 items in each; 10 out of 22 and 10 out of 22

OpenStudy (anonymous):

I've been getting problems like these on my h.w, but I get stuck on how to set up the problems as well, besides trying to figure out which formula to use

OpenStudy (amistre64):

if you spend all your life trying to find formulas, youll never learn the solutions :)

OpenStudy (anonymous):

im an independent student, there are tutors in my school but they don't help much.. :/

OpenStudy (amistre64):

what is the fundamental calculation for probability?

OpenStudy (anonymous):

uhmm, depends on events and outcomes right? dunno, this is supposedly geometry ..

OpenStudy (anonymous):

trying to do a quiz too but it wont let me go back to study the permutations/combinations

OpenStudy (amistre64):

the number of favorable outcomes, divided by the total number of possible outcomes

OpenStudy (amistre64):

it this case, we have 10 favorable oucomes ... since theres 10 candy bars there are 22 items altogether in the basket. the probabilty of pulling a candy is therefore: 10/22, and its the same for the next basket pull since it still has 10 out of 22

OpenStudy (amistre64):

now if you havent learned yet; when outcomes are independant, we can multiply thier probabilties so that gives us: 10/22 * 10/22 = (10/22)^2 for this case

OpenStudy (anonymous):

I see what you mean, that would be though it doesn't seem to be an answer, see, that would mean it would be 0.20, which would be 20/100 right?

OpenStudy (amistre64):

5/11)^2 = 25/121

OpenStudy (amistre64):

ive seen papers with typos many many times before. but unless if missed something, then there isnt much more to work with

OpenStudy (amistre64):

@rational any thoughts?

OpenStudy (anonymous):

how are the candy bars and chips distributed? does each basket have 10 candy bars and 12 chips or ?

OpenStudy (amistre64):

right, the information given is rather vague

OpenStudy (anonymous):

each basket has 12 bags of chips and 10 candy bars